# Deep Hedging with Options

### v2

@@ 4.5   Assessing the presence of speculative components in hedging positions
This section examines whether the RL risk management includes speculative elements, such as strategies that reap the time-varying risk premia embedded in hedging instruments. In what follows, the risk premium (RP) is defined as the difference between the discounted expected payoff and the option price at time $t$, i.e.,
@@
where $K^{*}$ is the hedging option strike price, the expectation is under the physical measure and $\mathcal{F}_{t}$ denotes the information available at time $t$ [^10] The usual definition of the risk premium is a return difference. However, when options are DOTM and their value is very low, this definition leads to numerical instability. The risk premium is estimated using a stochastic-on-stochastic simulation approach, where the present value of the expected payoff is computed through a nested simulation at each time step within the simulated paths.
This analysis investigates whether a statistical relationship exists between the risk premium $\mbox{RP}_{t}$ and the hedging position $\phi_{t+1}^{(\text{O})}$. Figure 4 presents a scatter plot of ranked data for these variables, using 20,000 samples [^11] A sample is a time point within a given path. from the 100,000 out-of-sample paths, which is repeated for each risk measure used in the optimization. The plot reveals no strong dependence patterns, suggesting a weak or insignificant relationship. This finding is further supported by sample correlations ranging from -0.001 to -0.006 across all risk measures, indicating that RL agents do not systematically seek to capture risk premium benefits.
@@
  • Results are computed using a sample of 20,000 data points from bcthe 100,000 out-of-sample paths. The hedged position is an ATM straddle with a maturity of $T=63$ days. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. Transaction cost levels are set to 0%.
As a complementary analysis, we examine whether our approach embeds speculative elements, such as statistical arbitrage overlays, that may deviate from sound risk management practices. Our results indicate that RL agents do not engage in such strategies, regardless of the risk measure used in optimization. Further details are provided in Appendix F.
@@ 4.6.1   Comparison with benchmarks
We analyze the relationship between the hedging option positions recommended by the DG strategy and those generated by RL agents. This analysis aims to understand how the RL outperformance documented in Sections 4.4.1 and 4.4.2 emerges by studying the positions taken by the hedger. Figure 5 presents the daily sample correlation between DG and RL hedging option positions, $\phi^{(\text{O},\text{DG})}_{t}$ and $\phi^{(\text{O},\text{RL})}_{t}$, under three risk measures: MSE, SMSE, and CVaR95%. The correlation is computed over the entire hedging period for two scenarios: one without transaction costs and another with $\kappa_{1}=0.05\%$ and $\kappa_{2}=1\%$ for illustration.
@@
  • Results are based on a sample of 100,000 out-of-sample paths. Agents are trained under the conditions described in Section 4.3. The hedged position is an ATM straddle with a maturity of $T=63$ day. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days.
Our numerical results reveal a consistent pattern across all risk measures, highlighting a significant divergence between RL and DG hedging strategies in terms of correlation, particularly at the start of the hedging horizon. Indeed, the RL agent benefits from learning experience to anticipate the future movements of state variables over multiple future periods. By contrast, the DG hedging agent is myopic in that he readjusts his hedging positions based on local risk. As time-to-maturity shrinks, both strategies become more similar. The inclusion of transaction costs leads the RL agent to maintain a distinct approach, with correlation remaining near zero for a significant portion of the hedging horizon. This is because the myopic DG agent does not have the same flexibility in managing transaction costs.
Additionally, a potential secondary source of divergence between these strategies may stem from differences in rebalancing size. While the frequency of rebalancing influences the timing of adjustments, the magnitude of these adjustments plays a key role in differentiating the hedging behaviors of the various strategies. Figure 6 illustrates the average hedging option position, along with the interquartile range, over time for all risk measures. The analysis is presented for two scenarios: one without transaction costs (first row), and another with transaction costs set to $\kappa_{1}=0.05\%$ and $\kappa_{2}=1\%$ (second row).
@@
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.1. The hedged position is an ATM straddle with a maturity of $T=63$ days. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. IQR stands for the interquartile range, representing the range between the 25th and 75th percentiles.
Our findings indicate that RL agents tend to hold smaller option positions during the early stages of the hedging period, a trend that is more pronounced with the introduction of transaction costs. This behavior arises from the substantial transaction cost associated with the hedging option, suggesting that RL agents favor more frequent rebalancing with smaller initial positions, gradually increasing their hedging positions over time. By deferring full engagement with the hedge, the RL agent seeks to balance cost efficiency with effective risk management, avoiding taking positions that might need to be unwound shortly after. Additionally, lower option positions in early stages allow the agent to initially limit the (short) exposure to the volatility risk premium while progressively scaling up the hedging positions. Thus, RL agents achieve twofold cost reductions, where both explicit transaction costs and implicit costs related to short exposure to the volatility risk premium are managed. In contrast, DG strategies adopt larger option positions early in the period to fully neutralize gamma risk. However, this approach leads to prolonged exposure to the volatility premium, making it suboptimal.
@@ 4.6.2   Sensitivity analysis
We analyze the sensitivity of RL agents’ positions to variations in the risk factors defining the IV surface, examining how they leverage information from its shape. Our analysis begins by evaluating RL policy behavior across different initial scenarios for the state variables ($(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$.
To assess the impact of each state variable, we sort the initial state vectors in the test set according to each variable and observe the corresponding hedging positions in the same order. This method accounts for the interdependence between these state variables and the broader state vector components, as detailed in Table 1, and reveals how changes in a selected variable influence hedging decisions. We focus on the initial state vector to ensure comparability across market conditions, particularly in terms of the initial underlying asset price and maturity, i.e., at $T=63$ days-to-maturity.
Figure 7 presents the hedging positions of the RL agent trained with the MSE risk measure under a no-transaction-cost scenario. Each panel displays the hedging positions when the initial state vectors are sorted according to each state variable, ($(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$.
@@
  • Results are computed using a sample of 20,000 data points from 100,000 out-of-sample paths for an ATM straddle with maturity of $T=63$ days. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. Transaction cost levels are set to 0%.
These empirical results suggest that the position in the hedging option exhibits a decreasing trend with respect to the conditional variance of the underlying asset returns, the long-term ATM level $\beta_{1}$ and the time-to-maturity slope $\beta_{2}$ of the IV surface. [^12] By contrast, there is no clear pattern related with the other factors as shown in panels C, D and E. As noted in François et al., (2024), this highlights that RL agents utilize both the historical variance process and market expectations of future volatility to adjust their positions. For instance, smaller positions on the hedging option when $\beta_{1}$, $\beta_{2}$ or $\sqrt{h_{R}}$ are higher can be explained by the higher cost of hedging in such circumstances. Indeed, both option prices and associated proportional transaction costs are higher.
@@ 4.7   Tracking error analysis
The differences between positions of RL and DG agents highlighted in previous sections allow RL agents to achieve higher performance with respect to terminal hedging error. This section investigates whether RL agents also retains good tracking performance before maturity.
We analyze the time-$t$ tracking error $\xi_{t}^{(\tilde{\phi}_{\theta},\,l)}$ defined in Equation (5) across all test set paths for different strategies throughout the hedging period. This comparison is conducted by evaluating three key metrics on each rebalancing day $t$: the average tracking error (ATE), root-mean squared tracking error (RMSTE), and semi root-mean squared tracking error (SRMSTE), given respectively by
$\text{ATE}=\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{(\tilde{\phi}_{\theta},\,l)},\,\text{RMSTE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{t,i}^{(\tilde{\phi}_{\theta},\,l)}\right)^{2}},\,\text{SRMSTE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{t,i}^{(\tilde{\phi}_{\theta},\,l)}\mathbbm{1}_{\left\{\xi_{t,i}^{(\tilde{\phi}_{\theta},\,l)}>0\right\}}\right)^{2}},$
where $\xi_{t,i}^{(\tilde{\phi}_{\theta},\,l)}$ represents the time-$t$ tracking error of the $i$-th path in the test set.
@@
  • Results are computed over 100,000 out-of-sample paths under the conditions outlined in Section 4.3.1. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days.
Figure 8 presents the evolution of these metrics over the hedging period under two scenarios: Panel A without transaction costs and Panel B with transaction costs. Panel B accounts for multiple DG strategies, each corresponding to a different optimal threshold $l$. The results indicate that, regardless of transaction costs, both the standard and asymmetric tracking error metrics (columns 2 and 3 of Figure 8) exhibit a monotonic upward trend for DG strategies. In contrast, RL strategies lead to curves that flatten out or even decrease through time demonstrating their ability to correct for past errors. Conversely, DG strategies are purely forward-looking, leading to the accumulation of unaddressed errors over time.
Furthermore, columns 2 and 3 show that RL agents maintain strong option-tracking performance in the absence of transaction costs, despite adopting strategies that differ from those derived using the DG approach. However, once transaction costs are introduced (panels B2 and B3 of Figure 8), the RL agent trained under the CVaR risk measure exhibits larger tracking error. This is primarily driven by the nature of the objective function, which focuses on minimizing the tail of losses only at the end of the hedging period. As a result, early deviations between the hedging and target portfolios do not necessarily lead to a loss in the tail of the distribution, and therefore do not require immediate correction, as positions can be rebalanced closer to maturity while keeping the CVaR at low levels.
@@
4.8   Backtesting
  • Results are computed based on the observed P&L from hedging 296 straddle positions with maturity 63-days under real market conditions observed from May 1, 1996, to December 31, 2020. A new ATM straddle is considered every 21 business days. Agents are trained according to the conditions outlined in Section 4.3 using an ATM call option with a maturity of $T^{*}=84$ days as the hedging instrument.
@@ 5   Conclusion
This study develops a deep hedging framework to manage the risk associated with S&P 500 options with a hedging portfolio including both options and underlying asset shares. In our work the information related to implied volatility surfaces is included within the set of state variables. The key differentiating aspect of our work is that with this information in hand, the adjustments in hedging positions not only integrate forward-looking expectations of market dynamics, but also capture the current price levels for options (and the associated volatility risk premium) within rebalancing decisions. The IV surface, conveniently represented by a parametric form, proves to be instrumental in refining the hedging policy.
Our hedging framework is also enhanced with a couple of trading features: (i) state-dependent no-trade regions to optimize rebalancing frequency in the presence of transaction costs, (ii) a soft constraint to mitigate speculative behavior, ensuring that hedging strategies focus on effective risk management.
Our approach consistently outperforms traditional benchmarks both with and without transaction costs, highlighting the hedging benefits of incorporating additional instruments, such as options. Furthermore, the inclusion of no-trade regions improves performance for both reinforcement learning and delta-gamma strategies: The former reduces unnecessary rebalancing, while the latter behaves like semi-static hedging approaches.
@@
Our study further documents the reasons driving the hedging outperformance of the reinforcement learning agent. In contrast to the myopic delta-gamma hedging, deep hedging begins with smaller option positions. This leads to less transaction costs and, more importantly, provides with more flexibility for appropriately rebalancing the hedging portfolio when uncertainty about the final moneyness of the position to hedge is gradually resolved. Smaller early-stage positions in the hedging option also reduce exposure to the volatility risk premium, leading to lower losses.
We show that reinforcement learning agents effectively incorporate both historical variance and market expectations of future volatility into their hedging decisions. The observed decline in hedging option positions in response to higher conditional variance, long-term ATM implied volatility level and time-to-maturity slope underscores the agents’ ability to dynamically mitigate risk, acting as a protective mechanism against volatility fluctuations.
@@
Backtests using historical data and various levels of transaction costs show that the reinforcement learning hedging performance is robust to diverse market conditions. They confirm that deep hedging with options using the implied volatility surface is a sound and practically applicable hedging approach.
Appendix A Neural network settings
A.1   Details for the MSGD training approach

### v3

@@ 4.5  Assessing the presence of speculative components in hedging positions
This section examines whether the RL risk management includes speculative elements, such as strategies that reap the time-varying risk premia embedded in hedging instruments. The risk premium (RP) is defined as the difference between the discounted expected payoff and the option price at time $t$, i.e.,
@@
where $K^{*}$ is the hedging option strike price, the expectation is under the physical measure and $\mathcal{F}_{t}$ denotes the information available at time tt. [^8] The usual definition of the risk premium is a return difference. However, when options are DOTM and their value is very low, this definition leads to numerical instability. The risk premium is estimated using a stochastic-on-stochastic simulation approach, where the present value of the expected payoff is computed through a nested simulation at each time step within the simulated paths.
@@
  • Results are computed using a sample of 20,000 data points from the 100,000 out-of-sample paths. The hedged position is an ATM straddle with a maturity of $T=63$ days. The hedging instrument is an ATM call option with an initial maturity of $T^{*}=84$ days. Transaction cost levels are set to 0%.
We investigate whether a statistical relationship exists between the risk premium $\mbox{RP}_{t}$ and the hedging position $\phi_{t+1}^{(\text{O})}$. Figure 5 presents a scatter plot of ranked data for these variables, using 20,000 samples from the 100,000 out-of-sample paths, which is repeated for the three risk measures. The plot reveals no strong dependence patterns, suggesting a weak or insignificant relationship. This finding is further supported by sample correlations ranging from -0.001 to -0.006, indicating that RL agents do not systematically seek to capture risk premium benefits.
As a complementary analysis, we examine whether our approach embeds speculative elements, such as statistical arbitrage overlays, that may deviate from sound risk management practices. Our results indicate that RL agents do not engage in such strategies, regardless of the risk measure used in optimization. Further details are provided in Appendix˜F.
@@ 4.6.1 Comparison with benchmarks
We analyze the relationship between the hedging option positions produced by the DG strategy and those generated by RL agents. This analysis aims to understand how the RL outperformance documented in Section˜4.4.1 and Section˜4.4.2 emerges by studying the positions taken by the hedger. Figure 6 presents for various days $t$, the sample correlation between DG and RL hedging option positions, $\phi^{(\text{O},\text{DG})}_{t}$ and $\phi^{(\text{O},\text{RL})}_{t}$, under the MSE, SMSE, and CVaR95% risk measures. The correlation is computed for two scenarios: one without transaction costs and another with $\kappa_{1}=0.05\%$ and $\kappa_{2}=1\%$ for illustration.
@@
  • Results are based on a sample of 100,000 out-of-sample paths. Agents are trained under the conditions described in Section˜4.3. The hedged position is an ATM straddle with a maturity of $T=63$ day. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days.
Our numerical results reveal a consistent pattern across all risk measures, highlighting a significant divergence between RL and DG hedging strategies in terms of correlation, particularly at the start of the hedging horizon. Indeed, the RL agent benefits from learning experience to anticipate the future movements of state variables over multiple future periods. By contrast, the DG hedging agent is myopic in that he readjusts his hedging positions based on local risk. As time-to-maturity shrinks, both strategies become more similar. The inclusion of transaction costs leads the RL agent to maintain a distinct approach, with correlation remaining near zero for a significant portion of the hedging horizon.
A potential secondary source of divergence between these strategies stems from differences in rebalancing size. While the rebalancing frequency influences the timing of adjustments, the magnitude of these adjustments plays a key role in differentiating the hedging behaviors. Figure 7 illustrates the average hedging option position, along with the interquartile range, over time for all risk measures. The analysis is presented for two scenarios: one without transaction costs (first row), and another with transaction costs set to $\kappa_{1}=0.05\%$ and $\kappa_{2}=1\%$ (second row).
@@
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section˜4.3.1. The hedged position is an ATM straddle with a maturity of $T=63$ days. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. IQR stands for the interquartile range, representing the range between the 25th and 75th percentiles.
Our findings indicate that RL agents tend to hold smaller option positions during the early stages of the hedging period, a trend that is more pronounced with the introduction of transaction costs. This behavior arises from the substantial transaction cost associated with the hedging option, suggesting that RL agents favor more frequent rebalancing with smaller initial positions, gradually increasing their hedging positions over time. By deferring full engagement with the hedge, the RL agent seeks to balance cost efficiency with effective risk management, avoiding taking positions that might need to be unwound shortly after. Additionally, lower option positions in early stages allow the agent to initially limit the (short) exposure to the variance risk premium while progressively scaling up the hedging positions. Thus, RL agents achieve twofold cost reductions, where both explicit transaction costs and implicit costs related to short exposure to the variance risk premium are managed. In contrast, DG strategies adopt larger option positions early in the period to fully neutralize gamma risk. However, this approach leads to prolonged exposure to the volatility premium, making it suboptimal.
@@ 4.6.2 Sensitivity analysis
We analyze the sensitivity of RL agents’ positions to variations in the risk factors defining the IV surface, examining how they leverage information from its shape. Our analysis begins by evaluating RL policy behavior across different initial scenarios for the state variables $(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$.
To assess the impact of each state variable, we sort the initial state vectors in the test set according to each variable and observe the corresponding hedging positions in the same order. This method accounts for the interdependence between these state variables and the broader state vector components, as detailed in Table 1, and reveals how changes in a selected variable influence hedging decisions.
Figure 8 presents the hedging positions of the RL agent trained with the MSE risk measure under a no-transaction-cost scenario. Each panel displays the hedging positions when the initial state vectors are sorted according to each state variable, $(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$.
These empirical results suggest that the position in the hedging option exhibits a decreasing trend with respect to the conditional variance of the underlying asset returns, the long-term ATM level $\beta_{1}$ and the time-to-maturity slope $\beta_{2}$ of the IV surface. [^9] By contrast, there is no clear pattern related with the other factors as shown in panels C, D and E. As noted in François et al., (2024), RL agents utilize both the historical variance process and market expectations of future volatility to adjust their positions. For instance, smaller positions on the hedging option when $\beta_{1}$, $\beta_{2}$ or $\sqrt{h_{R}}$ are higher can be explained by the higher cost of hedging in such circumstances. Indeed, both option prices and associated proportional transaction costs are higher.
@@
  • Results are computed using a sample of 20,000 data points from 100,000 out-of-sample paths for an ATM straddle with maturity of $T=63$ days. The hedging instrument is an ATM call option with an initial maturity of $T^{*}=84$ days. Transaction cost levels are set to 0%.
@@ 4.7  Tracking error analysis
The differences between positions of RL and DG agents allow RL agents to achieve higher performance with respect to terminal hedging error. This section investigates whether RL agents also retains good tracking performance before maturity.
We analyze the time-$t$ tracking error $\xi_{t}^{\tilde{\phi}_{\theta}}$ defined in Equation (5) across all test set paths throughout the hedging period. This comparison is conducted by evaluating three key metrics on each rebalancing day $t$: the average tracking error (ATE), root-mean squared tracking error (RMSTE), and semi root-mean squared tracking error (SRMSTE), given respectively by
$\text{ATE}=\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{(\tilde{\phi}_{\theta},\,l)},\,\text{RMSTE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{t,i}^{(\tilde{\phi}_{\theta})}\right)^{2}},\,\text{SRMSTE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{t,i}^{(\tilde{\phi}_{\theta})}\mathbbm{1}_{\left\{\xi_{t,i}^{(\tilde{\phi}_{\theta})}>0\right\}}\right)^{2}},$
where $\xi_{t,i}^{\tilde{\phi}_{\theta}}$ represents the time-$t$ tracking error of the $i$-th path in the test set.
@@
  • Results are computed over 100,000 out-of-sample paths under the conditions outlined in Section˜4.3.1. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days.
Figure 9 presents the evolution of these metrics over the hedging period under two scenarios: without transaction costs (Panel A) and with transaction costs (Panel B). Panel B accounts for multiple DG strategies, each corresponding to a different optimal no-trade threshold $\ell$. The results indicate that, regardless of transaction costs, both the standard and asymmetric tracking error metrics (columns 2 and 3 of Figure 9) exhibit a monotonic upward trend for DG strategies. In contrast, RL strategies lead to curves that flatten out or even decrease through time, demonstrating their ability to correct for past errors. Conversely, DG strategies are purely forward-looking, leading to the accumulation of unaddressed errors over time.
Furthermore, columns 2 and 3 show that RL agents maintain strong option-tracking performance in the absence of transaction costs, despite adopting strategies that differ from those derived using the DG approach. However, once transaction costs are introduced (panels B2 and B3 of Figure 9), the RL agent trained under the CVaR risk measure exhibits larger tracking error. This is primarily driven by the nature of the objective function, which focuses on minimizing the tail of losses only at the end of the hedging period. As a result, early deviations between the hedging and target portfolios do not necessarily lead to a loss in the tail of the distribution, and therefore do not require immediate correction, as positions can be rebalanced closer to maturity while keeping the CVaR at low levels. Larger tracking errors in early stages are expected because the RL optimization leads to smaller hedging positions, see Figure 7.
@@
5  Out-of-sample backtesting
We assess the performance of our framework under actual market conditions, using historical option prices sourced from OptionMetrics observed between December 31, 2020, and October 31, 2023. We evaluate the hedging performance across 4,134 near-the-money 63-day European straddles, where the option strike lies within $\pm 10\%$ of the underlying asset’s initial price. Each straddle is hedged using a combination of a call option with a longer maturity (between 78 and 84 days of maturity depending on availability), the underlying asset and the cash account. [^10] On each day of the out-of-sample dataset, the IV parameters $\beta_{t}$ are estimated. We then re-estimate their joint dynamics with the S&P 500 returns to recreate the state space. To ensure consistency with the simulation environment used during training, all underlying price paths are rescaled to start at a normalized value of 100.
We compare the performance of the practitioners’ delta-gamma hedging with that of the RL algorithms with and without IV surface information. [^11] The RL benchmark without IV surface information is analogous to the proposed RL method, except that predictors $\beta_{t,1},\ldots,\beta_{t,5},h_{t,1},\ldots,h_{t,5}$ are dropped from the state space. Figure 10 shows that, in terms of MSE, the RL algorithm without IV surface information performs worse than the two other approaches. Interestingly, in the absence of transaction costs, the practitioners’ delta-gamma hedging and the RL algorithm with IV information exhibit very similar MSE. The RL-algorithm with the complete information slightly outperforms the other approaches in presence of transaction costs, again in terms of MSE. The main conclusion is that RL approaches do not necessarily dominate traditional methods; their performance critically depends on the information provided to the algorithm. However, in terms of tail risk, the RL algorithms clearly outperform the practitioners’ delta-gamma approach. Moreover, receiving information about the IV surface clearly improves the performance the RL algorithm.
Figure 10: Out-of-sample backtest performance metrics on hedging errors, with and without transaction costs.
  • The backtest is conducted on 4,134 around-the-money straddle intruments, using actual market prices observed between December 31, 2020 and October 31, 2023.
Further evidence of the RL approach’s superior performance is provided by the distribution of hedging errors (without transaction costs) shown in Figure 11. The first row compares the RL algorithm with the full information to the practitioners’ delta-gamma strategy. It is clear that the distribution of the practitioners’ delta-gamma strategy is shifted to the right and exhibits a heavier right tail. In the second row, we observe the benefits of incorporating IV information into the RL strategies. The distribution of hedging errors is more concentrated around zero when such information is included.
Figure 11: Distribution of hedging errors for near-at-the-money straddles.
  • The backtest is conducted on 4,134 around-the-money straddle intruments, using actual market prices observed between December 31, 2020 and October 31, 2023. No transaction costs are applied.
Figure 12: Time series of hedging strategies P&L and market conditions.
  • The backtest is conducted on 4,134 around-the-money straddle intruments, using actual market prices observed between December 31, 2020 and October 31, 2023. No transaction costs are applied.
Although risk management is primarily associated with measures of dispersion and downside risk, it is interesting to note that only the practitioners’ delta-gamma strategy exhibits a negative cumulative P&L (see Figure 12). Moreover, having the full information about the IV surface helps the RL algorithm achieve the best cumulative P&L. We observe that the cumulative P&L of the RL algorithm with the full information increases significantly in the second half of the sample. Examining market conditions (Panel B to Panel E), we see that this period is associated with IV slopes that are strongly positive (see Panel D). In the middle of the sample, there is a period of high volatility (see Panel E), but this information is captured by both the practitioners’ delta-gamma hedging and the RL algorithms.
These findings demonstrate that RL agents achieve consistent and competitive performance when applied to unseen historical market conditions, despite being trained on simulated data. Their ability to adapt to diverse environments and maintain superior risk control highlights the practical value of this approach in hedging tasks.
@@ 6  Conclusion
This study develops a deep hedging framework to manage the risk associated with S&P 500 options with a hedging portfolio including both options and underlying asset shares. In our work the information related to implied volatility surfaces is included within the set of state variables. The key differentiating aspect of our work is that with this information in hand, the adjustments in hedging positions not only integrate forward-looking expectations of market dynamics, but also capture the current price levels for options (and the associated variance risk premium) within rebalancing decisions. The IV surface, conveniently represented by a parametric form, proves to be instrumental in refining the hedging policy. A soft constraint is included in the optimization scheme to mitigate speculative behavior, ensuring that hedging strategies focus on effective risk management.
@@
Our approach consistently outperforms traditional benchmarks both with and without transaction costs. It also highlights the substantial hedging benefits of incorporating additional instruments, such as options. Our study further documents the reasons driving the hedging outperformance of the reinforcement learning agent. In contrast to the myopic delta-gamma hedging, deep hedging begins with smaller option positions. This leads to less transaction costs and, more importantly, provides more flexibility for appropriately rebalancing the hedging portfolio when uncertainty about the final moneyness of the position to hedge is gradually resolved. Smaller early-stage positions in the hedging option also reduce exposure to the variance risk premium, leading to lower losses. We show that reinforcement learning agents effectively incorporate both historical variance and market expectations of future volatility into their hedging decisions. The observed decline in hedging option positions in response to higher conditional variance, long-term ATM implied volatility level and time-to-maturity slope underscores the agents’ ability to dynamically mitigate risk, acting as a protective mechanism against volatility fluctuations.
@@
Out-of-sample backtests using historical data and various levels of transaction costs show that the reinforcement learning hedging performance is robust to diverse market conditions and superior to that of benchmarks in terms of downside risk management, on top of providing superior profitability. Such tests highlight the importance of information embedded in implied volatility surfaces. This confirms that deep hedging with options using the implied volatility surface is a sound and practically applicable hedging approach.
Appendices
Appendix A No trade region
At time $t$, the no-trade region121212 No-trade regions, which mitigate the impact of transaction costs, have been extensively studied in the portfolio optimization literature. Constantinides, (1986) first introduced the idea that proportional transaction costs give rise to such regions—a concept further developed by Davis and Norman, (1990) and Balduzzi and Lynch, (1999), who emphasized portfolio allocation over rebalancing costs. In the hedging context, optimal rebalancing based on delta variations has been explored by Henrotte, (1993), Toft, (1996), and Martellini and Priaulet, (2002). Hodges and Neuberger, (1989) and Clewlow and Hodges, (1997) examine hedging within a utility-maximization framework. The optimal hedging strategy consists of no-trade bands around delta, whose width depends on the hedger’s risk aversion. is determined by the distance between the current portfolio position, $\phi_{t}$, and the next position proposed by the ANN, $\tilde{\phi}_{\theta}(X_{t})$. Specifically, rebalancing occurs only if the cumulative deviation in positions across hedging instruments exceeds a threshold $l$:
$(\phi^{(S)}_{t+1},\phi^{(O)}_{t+1})$ $=\begin{cases}(\phi^{(S)}_{t},\phi^{(O)}_{t}),&\text{if }\lvert\phi_{t}^{(S)}-\tilde{\phi}_{\theta}^{(S)}(X_{t})\rvert+\lvert\phi^{(\text{O})}_{t}-\tilde{\phi}_{\theta}^{(\text{O})}(X_{t})\rvert\leq\ell,\\[8.0pt] \left(\tilde{\phi}^{(S)}_{\theta}(X_{t}),\,\tilde{\phi}^{(O)}_{\theta}(X_{t})\right),&\text{otherwise}.\end{cases}$ (10)
The bank account position is determined by the self-financing constraint (1). This formulation expresses the no-trade region in terms of the number of shares of option contracts, providing a measure of the distance at which rebalancing becomes cost-effective, capturing the trade-off between transaction costs and maintaining proximity to the desired portfolio adjustments. Indeed, when rebalancing actions proposed by the neural network are minor, they are not implemented because (i) this only leads to a small misalignment with the ideal hedging positions and (ii) this allows avoiding transaction costs. [^13] We tried other specifications for the no-trade region (for instance explicitly capturing transaction cost amounts), with results being qualitatively similar. The rebalancing threshold $\ell$ is treated as a learnable parameter included in the ANN parameters $\theta$, allowing the model to jointly optimize the size of rebalancing actions and decisions of whether or not to rebalance.
This analysis incorporates the no-trade region, defined by Equation (10), to optimize rebalancing frequency while accounting for transaction costs. For benchmarks, the rebalancing threshold $\ell$ is estimated using the approach described in Section˜D.3. In contrast, RL strategies estimate this parameter jointly with other ANN parameters during training.
Figure 13: Optimal rebalancing threshold $\ell$ values for DG and RL strategies.
Figure 13 reports the optimal rebalancing thresholds $\ell$ across different transaction cost levels for both DG (gray) and RL (blue) strategies, considering all risk measures. Remarkably, the RL algorithms barely rely on the no-trade region, as the optimal values of $\ell$ are close to zero. Figure 3 provides further evidence of this phenomenon: risk metrics associated with the RL approaches are minimally affected by the presence of the no-trade region. Such no-trade region is primarily introduced to assist the delta-gamma benchmarks, which are not inherently designed to handle transaction costs efficiently. In the absence of transaction costs, the optimal no-trade region parameter $\ell$ collapses to zero.
Appendix B Details for the MSGD training approach